# A Network-Based Measure of Cosponsorship Influence on Bill Passing in the United States House of Representatives

### v1

@@ Abstract
Political Networks, United States Congress, Cosponsorship Networks
@@ 1 Introduction
Since 1789, members of the United States Congress have been entrusted to help determine the laws that shape the US’s political landscape [1]. Congress has two chambers: the House of Representatives, which has 435 legislators (which are called representatives); and the Senate, which has 100 legislators (which are called senators). Congress decides which pieces of legislation to draft, which bills to vote on, and ultimately which bills to pass to present to the US President. Congress considers thousands of legislative proposals each year, but only a few hundred of them ultimately pass and are enacted into law [2]. Why do some bills pass but other seemingly comparably important bills fail?
@@
Political scientists have long considered the role that social networks play in politics [3, 4, 5, 6, 7, 8]. Humans are social creatures, and the choices that they make are influenced by those around them. There is evidence that interpersonal relationships influence political behavior, including in voter decisions during presidential elections [9, 10]. Naturally, one also expects the behavior of politicians themselves to be subject to influence from their peers. In both the House of Representatives and the Senate, researchers have observed that party affiliation influences roll-call votes beyond legislators’ political preferences [11, 12]. Accordingly, legislators are affected by the opinions of others in their social networks during their political decision-making processes.
@@
During the last two decades, many researchers have used network analysis to study social and political ties between legislators [13, 14, 15, 16]. Network analysis is an informative approach to examine the mutual dependencies and channels of influence between interacting individuals, groups, and institutions [17, 18]. It can thereby yield important insights into political processes, which involve complex relationships between many entities. Researchers have analyzed many types of political and social networks. These include joint membership of committees and subcommittees [19, 20], digital-trace data on social media [21], coappearances in World Wide Web searches [22], and bill cosponsorship networks [23, 24].
@@
An important application of networks in political science is the analysis of how groups of individuals debate, decide, and commit to decisions (such as reaching a verdict as a jury, allocating funds as a committee, or a passing a bill in Congress [15]). The analysis of bill cosponsorship networks can yield insight into the effects of relationships and networks on political processes [23, 24]. A proposed bill in Congress has one sponsor (the main member of Congress who introduces the bill) and can have any number (even a very large number) of cosponsors. Cosponsorship data captures a notion of collaboration, and it can also indicate information about the time evolution of relationships and social-tie strengths between Congress members. Many researchers have analyzed cosponsorship networks to obtain insights into political dynamics [23, 24, 25, 26, 27, 28]. For example, Fowler [23] formulated and analyzed a ‘connectedness’ centrality measure to identify important members of Congress in bill cosponsorship networks. This connectedness measure considers a weighted count — using both the numbers of cosponsors on bills and the numbers of bills that pairs of legislators both cosponsor — of bill cosponsorship between legislators, takes the reciprocal of this weight to yield a cosponsorship distance between each two legislators, and uses Dijkstra’s algorithm to find the shortest distance between each pair of legislators. Fowler then calculated a notion of connectedness for each legislator using their mean distance to the other legislators. Fowler illustrated that such connectedness can help successfully infer the passage of amendments to existing legislation, thereby using properties of bill cosponsorship networks as a proxy for legislator influence.
@@
In the present paper, we define a notion of ‘influence’ between Congress members that depends on achieving some level of legislative success. To examine such legislator influence, we consider bill cosponsorship relationships between members of the US House of Representatives. The incorporation of legislative outcomes into our notion of influence is a key distinction between our approach and the connectedness measure in [23]. We find that our influence is a better predictor of bill passage than the conventional network centrality measures of eigenvector centrality, closeness centrality, and strength centrality (i.e., weighted degree centrality).
@@
Our paper proceeds as follows. In Section 2, we contextualize bill cosponsorship in the US Congress and detail its relevance as a potential indicator of legislator influence. In Section 3, we discuss the data (including our preprocessing steps) that we employ in our investigation. In Section 4, we calculate influence in the US House of Representatives using bill cosponsorship data from January 2009 through January 2019 (i.e., the 111th–115th Congresses). In Section 5, we compare our notion of influence to eigenvector centrality. We conclude in Section 6 by discussing our results and outlining further research directions. In Appendix A, we compare our notion of influence to closeness centrality and strength centrality.
2 Cosponsorship History and Relevance to Legislator Influence
@@
Although bill cosponsorship is a common practice in the US Congress, scholars and pundits debate its significance in the legislative process [29]. Some researchers have argued that cosponsorship is a low-risk way for Congress members to signal to their constituents that they are following through on their campaign promises [29]. If a bill that a Congress member cosponsors passes, they can use it as evidence to convey to their constituents that they are doing good work. If a bill that they cosponsor does not pass, which is common in Congress, there are no real consequences. However, researchers have also found that the sheer number of cosponsors on a bill is not a strong indicator of its success [29], so it may be that bill cosponsorship does not play a significant role in the legislative process. A more nuanced perspective is that cosponsorship ties can indicate the evolving political relationships between members of Congress, which ultimately may play a role in legislation, even if cosponsor count itself does not give a clear signal of a bill’s success [23].
@@
Making decisions on which bills to cosponsor is a complex and nuanced process, and a Congress member may cosponsor a bill for a variety of reasons [31]. For example, they can support a bill because they believe that it is genuinely beneficial, because it appeases their constituents and may thus benefit their chances of reelection, or because it helps it helps them curry favor with other legislators. In the 111th–115th Congresses (i.e., between January 2009 and January 2019), which is the data that we study (see Section 3), the maximum percent of all proposed bills that were sponsored by an individual Congress member in a single Congress was 12.15% in the House of Representatives and 13.67% in the Senate. Because support from other people is important in Congress, it seems likely that legislator decision-making is affected by a desire to make or maintain personal relationships, which may provide social capital for the future. Accordingly, a Congress member may be more likely to cosponsor a bill with an individual with whom they have a successful relationship through bill passage [32]. Therefore, repeated bill cosponsorship between two or more Congress members may give evidence of legislative alliances, with patterns of supporting each other’s legislation [33]. Such quid pro quo is sometimes called “logrolling” in legislative and other organizational contexts [34]. There is empirical evidence for such support patterns, with the structure of some social networks in Congress resembling the network structure in the mutual licking of cows [35]. Legislator influence can also arise through homophily [26]. For example, collaborations within a party occur not only from mere party loyalty (and the ensuing promotion of a party’s agenda and “brand name”) but also from ideological similarities [36]. Some legislators also cosponsor bills across party lines (especially for Senators, where it is helpful to advance career goals [37]), even amidst growing partisanship in Congress [38]. Drawing from these ideas, we argue that the passage of a large percentage of bills that a Congress member supports (as demonstrated through cosponsorship) may be an indication that that Congress member is influential.
@@ 3 Data and Preprocessing
We extract cosponsorship activity on proposed legislation in the US House of Representatives between January 2009 and January 2019 (i.e., the 111th–115th Congresses). Congressional elections occur every two years, so our data spans five Congresses. This data is publicly available through ProPublica, who retrieved the data from the US Congress [39]. To match Congress members who served multiple terms, we leverage other data sets to complete missing identification numbers in our main cosponsorship data set [40, 41, 42]. In our analysis, we only consider bills; we exclude all other types of legislation (namely, simple resolutions, concurrent resolutions, and joint resolutions). Bills undergo a different legislative process than other types of legislation, so it is difficult to directly compare their success to the success of other types of legislation [43, 44]. Bills typically propose significant pieces of legislation; if successful, they carry the force of law, which is not the case of simple and concurrent resolutions. Joint resolutions are often used for continuing or emergency appropriations, so Congressional relationships are likely to have a different relevance for them than for bills. Bills constitute about 70% of all proposed legislation in Congresses 111–115 [39], so using only bills leaves us with ample data to analyze.
In Figure 1, we illustrate the life cycle of a proposed bill [45]. A bill is drafted in either the Senate or the House of Representatives. If a bill passes in its originating chamber of Congress, it is sent to the other chamber for review. In our investigation, we only examine whether or not bills pass in their original chamber. We study bills in the US House of Representatives, and we say that a bill “passes” if it succeeds at the introduction, committee, and debate stages in Figure 1 (irrespective of whether or not it passes in the Senate and whether or not it is enacted into law). We say that a bill “fails” if it fails at some stage in the House. We limit our analysis in this way for two primary reasons. First, Congress members can only cosponsor bills that are introduced in their own chamber [45]. Therefore, it seems reasonable to study the effects of cosponsorship within a single Congressional chamber. Second, very few bills survive the legislative process to be enacted into law. Roughly 3% of all bills in the ProPublica data (which covers both the House and the Senate) were ultimately enacted into law [39].
@@
There are differences in cosponsorship behaviors in the Senate and the House of Representatives [46]. The House of Representatives (with 435 representatives) is much larger than the Senate (which has 100 senators). Therefore, the House seemingly has more opportunities for cosponsorships and perhaps has a greater incentive to form political relationships. Representatives serve two-year terms, whereas senators serve six-year terms [45]. Shorter terms entail more frequent reelection pressures and hence can hold individuals more accountable to their constituents. Unsurprisingly, cosponsorship activity increases dramatically in election years [23]. Additionally, because there are more legislators in the House, it may be more likely for a legislator to find a group of individuals with whom they align well politically.
In Table 1, we summarize key statistics from our data set [39] on bill passage in the House of Representatives.
@@
Table 1: Summary of data for bill passage in the House of Representatives in Congresses 111–115. In the first row, we show the percent of bills originating in the House of Representatives that pass in the House. In the second row, we show the percent of bills originating in the House of Representatives that were enacted into law. In the third row, we show the mean number of cosponsors of a bill. In the fourth row, we show the maximum number of cosponsors of a bill. In the fifth row, we show the mean number of bills that a representative cosponsors. In the sixth row, we show the maximum number of bills that a representative cosponsors.
@@ 4 Defining Legislator Influence
To construct our influence score, we consider the influence that each representative has with all other representatives at some time $t$. We examine such influence using a bill cosponsorship network in which each representative is a node and an undirected edge exists between representatives $i$ and $j$ if they sponsor or cosponsor a bill. We do not include any self-edges. We treat sponsorship and cosponsorship of bills on equal footing, rather than considering directed edges from bill cosponsors to sponsors. For simplicity, we use the term “cosponsor” to refer to any individual that either sponsors or cosponsors a bill. Our cosponsorship network is weighted, with time-dependent edge weights. The weight of an edge at time $t$ depends on the number of bills that two representatives both cosponsor at all times up to and including $t$, with decay as cosponsorship recedes into the past. In Section 4, we give a detailed discussion of how we determine edge weights.
To obtain a tensor representation of our bill cosponsorship network, we first construct an adjacency tensor with entries $A[t,i,j]_{\text{pass}}$ that encode the number of bills that include both representatives $i$ and $j$ as cosponsors, are introduced in month $t$, and eventually pass (including possibly in month $t$) in the House of Representatives. We also construct an adjacency tensor with entries $A[t,i,j]_{\text{tot}}$ that encodes the total number of bills, irrespective of their success, that include both representatives $i$ and $j$ as cosponsors and are introduced in month $t$. We suppose that representatives accumulate influence together when the bills that they mutually cosponsor pass. Accordingly, we assume that the influence between representatives $i$ and $j$ is symmetric. Therefore, $A[t,i,j]_{\text{pass}}$ = $A[t,j,i]_{\text{pass}}$ and $A[t,i,j]_{\text{tot}}=A[t,j,i]_{\text{tot}}$. It future work, it is desirable to also consider asymmetric influence between sponsors and cosponsors.
@@ 4.1 Accounting for the Accumulation and Decay of Influence with Time
We seek a cumulative-in-time notion of influence that incorporates decay with time. Intuitively, a legislator that has been in Congress for ten years has more ‘influence capital’ (i.e., social capital for the purpose of influence) than a legislator on their first day in Congress [47]. Additionally, a legislator that was very influential many years ago who has been inactive for a long time likely has experienced decay in their influence capital. To account for both of these ideas, we use an exponential decay function $e^{kt}$ so that accrued influence (from successful bills) decays as a function of the time $t$ since the introduction of a bill. The decay rate $k$ controls the half-life of a unit of influence between individuals that have successfully cosponsored a single passed bill. We examine influence half-lives of 6, 12, and 24 months by setting $k=\frac{\text{ln}(0.5)}{\text{half-life}}{\approx}\frac{-0.693}{\text{half-life}}$.
We consider an exponential decay of influence for both $A[t,i,j]_{\text{pass}}$ (the number of bills that are introduced in the House at time $t$ that include both representatives $i$ and $j$ as cosponsors and eventually pass in the House) and $A[t,i,j]_{\text{tot}}$ (the number of bills that are introduced in the House at time $t$, irrespective of their success, that include both representatives $i$ and $j$). We then aggregate the decayed results month by month to yield tensors $C_{\text{pass}}$ and $C_{\text{tot}}$ with entries
@@
The entry $C[t,i,j]_{\text{pass}}$ is the number of bills that are cosponsored by both representatives $i$ and $j$ and pass in the House between months $1$ and $t$ (inclusive of months $1$ and $t$), with a separate decay applied at each month. For example, we obtain the influence between representatives $i$ and $j$ at month 3 (while accounting for the influence that they accumulate in months $1$ and $2$) by calculating
@@
which incorporates a decay at each month. Analogously, the entry $C[t,i,j]_{\text{tot}}$ is the total number of bills (irrespective of their success) that are cosponsored by both representatives $i$ and $j$ between months $1$ and $t$, with decay at each month.

### v2

# A Network-Based Measure of Cosponsorship Influence on Bill Passage in the United States House of Representatives

@@ Abstract
Political Networks, United States Congress, United States House of Representatives, Cosponsorship Networks, legislative influence, centrality
@@ 1 Introduction
Since 1789, members of the United States Congress have been entrusted to help determine the laws that shape the US’s political landscape [1]. Congress has two chambers: the House of Representatives, which has 435 legislators (which are called representatives); and the Senate, which has 100 legislators (which are called senators). Congress decides which pieces of legislation to draft, which bills to vote on, and ultimately which bills to pass to present to the US President. Congress considers thousands of legislative proposals each year, but only a few hundred of them ultimately pass and are enacted into law [2]. Why do some bills pass but other bills fail?
@@
Political scientists have long considered the role that social networks play in politics [3, 4, 5, 6, 7, 8]. Humans are social creatures, and the choices that they make are influenced by the people and environments around them. There is evidence that interpersonal relationships influence political behavior, including in voter decisions during presidential elections [9, 10]. Naturally, one also expects the behavior of politicians themselves to be subject to influence from their peers. In both the House of Representatives and the Senate, researchers have observed that party affiliation influences roll-call votes beyond legislators’ political preferences [11, 12]. Accordingly, legislators are affected by the opinions of others in their social networks during their political decision-making processes.
@@
During the last two decades, many researchers have used network analysis to study social and political ties between legislators [13, 14, 15, 16]. More generally, network analysis is an informative approach to examine the mutual dependencies and channels of influence between interacting individuals, groups, and institutions [17, 18]. It can thus yield important insights into political processes, which involve complex relationships between many entities. Researchers have analyzed many types of political and social networks. These include joint membership of legislative committees and subcommittees [19, 20], digital-trace data on social media [21], coappearances in World Wide Web searches [22], and bill cosponsorship networks [23, 24].
@@
An important application of networks in political science is the analysis of how groups of individuals debate, decide, and commit to decisions (such as reaching a verdict as a jury, allocating funds as a committee, or a passing a bill in Congress [15]). The analysis of bill cosponsorship networks can yield insight into the effects of relationships and networks on political processes [23, 24]. A proposed bill in Congress has one sponsor (the main member of Congress that introduces the bill) and can have any number (even a very large number) of cosponsors. Cosponsorship data captures a notion of collaboration, and it can also indicate information about the time evolution of relationships and social-tie strengths between Congress members. Many researchers have analyzed cosponsorship networks to obtain insights into political dynamics [23, 24, 25, 26, 27, 28]. For example, Fowler [23] formulated and analyzed a ‘connectedness’ centrality measure to identify important members of Congress in bill cosponsorship networks. This connectedness measure considers a weighted count — using both the numbers of cosponsors on bills and the numbers of bills that pairs of legislators both cosponsor — of bill cosponsorship between legislators, takes the reciprocal of this weight to yield a cosponsorship distance between each two legislators, and uses Dijkstra’s algorithm to find the shortest distance between each pair of legislators. Fowler then calculated a notion of connectedness for each legislator using that legislaor’s mean distance to the other legislators. Fowler illustrated that such connectedness can help successfully infer the passage of amendments to existing legislation, thereby using properties of bill cosponsorship networks as a proxy for legislator influence.
@@
In the present paper, we define a notion of ‘influence’ between Congress members that depends on achieving some level of legislative success. To examine such legislator influence, we consider bill cosponsorship relationships between members of the US House of Representatives. The incorporation of legislative outcomes into our notion of influence is a key distinction between our approach and the connectedness measure in [23]. We find that our notion of influence is a better predictor of bill passage than several conventional network centrality measures (specifically, eigenvector centrality, closeness centrality, and strength centrality).
@@
Our paper proceeds as follows. In Section 2, we contextualize bill cosponsorship in the US Congress and detail its relevance as a potential indicator of legislator influence. In Section 3, we discuss the data (including our preprocessing steps) that we employ in our investigation. In Section 4, we calculate influence in the US House of Representatives using bill cosponsorship data from January 2009 through January 2019 (i.e., the 111th–115th Congresses). In Section 5, we compare our notion of influence to eigenvector centrality. We conclude in Section 6 by discussing our results and outlining further research directions. In Appendix A, we compare our notion of influence to closeness centrality and strength centrality. Our code is publicly available at https://github.com/sarahsotoudeh/LegislativeInfluence.
2 Cosponsorship History and Relevance of Cosponsorship to Legislator Influence
@@
Although bill cosponsorship is a common practice in the US Congress, scholars and pundits debate its significance in the legislative process [29]. Some researchers have argued that cosponsorship is a low-risk way for Congress members to signal to their constituents that they are following through on their campaign promises [29]. If a bill that a Congress member cosponsors passes, they can use it as evidence to convey to their constituents that they are doing good work. If a bill that they cosponsor does not pass, which is common in Congress, there are no real consequences. Researchers have also found that the sheer number of cosponsors on a bill is not a strong indicator of its success [29], so perhaps bill cosponsorship does not play a significant role in the legislative process. A more nuanced perspective is that cosponsorship ties can indicate the evolving political relationships between members of Congress, which ultimately may play a role in legislation, even if cosponsor count itself does not give a clear signal of a bill’s success [23].
Making decisions on which bills to cosponsor is a complex and nuanced process, and a Congress member may cosponsor a bill for a variety of reasons [31]. For example, they can support a bill because they believe that it is genuinely beneficial, because it appeases their constituents and may thus benefit their chances of reelection, or because it helps them curry favor with other legislators. In the 111th–115th Congresses (i.e., between January 2009 and January 2019), which is the data that we study (see Section 3), the overwhelming majority of bills had at least one cosponsor. In the 111th–115th Congresses, at least 87.85% of the bills in the House of Representatives had cosponsors in each of the 10 terms during these 5 Congresses. During the same period, at least 86.33% of the bills in the Senate had cosponsors in each of the 10 terms.
@@
Because support from other people is important in Congress, it seems likely that legislator decision-making is affected by a desire to make or maintain personal relationships, which may provide social capital for the future. Accordingly, a Congress member may be more likely to cosponsor a bill with an individual with whom they have a successful relationship through bill passage [32]. Therefore, repeated bill cosponsorship between two or more Congress members may give evidence of legislative alliances, with patterns of supporting each other’s legislation [33]. Such quid pro quo is sometimes called ‘logrolling’ in legislative and other organizational contexts [34]. There is empirical evidence for such support patterns, with the structure of some social networks in Congress resembling the network structure in the mutual licking of cows [35]. Legislator influence can also arise through homophily [26]. For example, collaborations within a party occur not only from mere party loyalty (and the ensuing promotion of a party’s agenda and ‘brand name’) but also from ideological similarities [36]. Some legislators also cosponsor bills across party lines (especially for Senators, where it is helpful to advance career goals [37]), even amidst growing partisanship in Congress [38]. Drawing from these ideas, we argue that the passage of a large percentage of bills that a Congress member supports (as demonstrated through cosponsorship) may be an indication that that Congress member is influential.
@@ 3 Data and Preprocessing
We extract cosponsorship activity on proposed legislation in the US House of Representatives between January 2009 and January 2019 (i.e., the 111th–115th Congresses). Congressional elections occur every two years, so our data spans five Congresses. This data is publicly available through ProPublica, which retrieved the data from the US Congress [39]. To match Congress members who served multiple terms, we leverage other data sets to complete missing identification numbers in our main cosponsorship data set [40, 41, 42]. In our analysis, we only consider bills; we exclude all other types of legislation (namely, simple resolutions, concurrent resolutions, and joint resolutions). Bills undergo a different legislative process than other types of legislation, so it is difficult to directly compare their success to the success of other types of legislation [43, 44]. Bills typically propose significant pieces of legislation; if successful, they carry the force of law, which is not the case of simple and concurrent resolutions. Joint resolutions are often used for continuing or emergency appropriations, so Congressional relationships are likely to have a different relevance for them than for bills. Bills constitute about 70% of all proposed legislation in Congresses 111–115 [39], so using only bills leaves us with ample data to analyze.
In Figure 1, we illustrate the life cycle of a proposed bill [45]. A bill is drafted in either the Senate or the House of Representatives. If a bill passes in its originating chamber of Congress, it is sent to the other chamber for review. In our investigation, we only examine whether or not bills pass in their original chamber. We study bills in the US House of Representatives, and we say that a bill ‘passes’ if it succeeds at the introduction, committee, and debate stages in Figure 1 (irrespective of whether or not it passes in the Senate and of whether or not it is enacted into law). We say that a bill ‘fails’ if it fails at some stage in the House. We limit our analysis in this way for two primary reasons. First, Congress members can only cosponsor bills that are introduced in their own chamber [45]. Therefore, it seems reasonable to study the effects of cosponsorship within a single Congressional chamber. Second, very few bills survive the legislative process to be enacted into law. Roughly 3% of all bills in Congresses 111–115 in the ProPublica data (which covers both the House and the Senate) were ultimately enacted into law [39].
@@
There are differences in cosponsorship behaviors in the Senate and the House of Representatives [46]. The House of Representatives (which has 435 representatives) is much larger than the Senate (which has 100 senators). Therefore, the House seemingly has more opportunities for cosponsorships and perhaps has a greater incentive to form political relationships. Representatives serve two-year terms, whereas senators serve six-year terms [45]. Shorter terms entail more frequent reelection pressures and hence can hold individuals more accountable to their constituents. Unsurprisingly, cosponsorship activity increases dramatically in election years [23]. Additionally, because there are more legislators in the House than in the Senate, it may be more likely for a legislator in the former chamber to find a group of individuals with whom they align well politically.
In Table 1, we summarize key statistics from the employed data set [39] on bill passage in the House of Representatives.
@@
Table 1: Summary of data for bill passage in the US House of Representatives in Congresses 111–115. In the first row, we show the percent of bills originating in the House of Representatives that pass in the House. In the second row, we show the percent of bills originating in the House of Representatives that were enacted into law. In the third row, we show the mean number of cosponsors of a bill. In the fourth row, we show the maximum number of cosponsors of a bill. In the fifth row, we show the mean number of bills that a representative cosponsors. In the sixth row, we show the maximum number of bills that a representative cosponsors.
@@ 4 Defining Legislator Influence
To construct our influence score, we consider the influence that each representative has both on and from all other representatives at some time $t$. We examine such influence using a bill cosponsorship network in which each representative is a node and an undirected edge exists between representatives $i$ and $j$ if they sponsor or cosponsor a bill. We do not include any self-edges. We treat sponsorship and cosponsorship of bills on equal footing, rather than considering directed edges from bill cosponsors to sponsors. For simplicity, we use the term ‘cosponsor’ to refer to any individual that either sponsors or cosponsors a bill. Our cosponsorship network is weighted, with time-dependent edge weights. The weight of an edge at time $t$ depends on the number of bills that two representatives both cosponsor at all times up to and including $t$, with decay as cosponsorship recedes into the past. In Section 4, we give a detailed discussion of how we determine edge weights.
To obtain a tensor representation of our bill cosponsorship network, we first construct an adjacency tensor with entries $A[t,i,j]_{\text{pass}}$ that encode the number of bills that include both representatives $i$ and $j$ as cosponsors, are introduced in month $t$, and eventually pass (including possibly in month $t$) in the House of Representatives. We also construct an adjacency tensor with entries $A[t,i,j]_{\text{tot}}$ that encodes the total number of bills, irrespective of their success, that include both representatives $i$ and $j$ as cosponsors and are introduced in month $t$. We suppose that representatives accumulate influence together when the bills that they mutually cosponsor pass. Accordingly, we assume that the influence between representatives $i$ and $j$ is symmetric. Therefore, $A[t,i,j]_{\text{pass}}$ = $A[t,j,i]_{\text{pass}}$ and $A[t,i,j]_{\text{tot}}=A[t,j,i]_{\text{tot}}$. In future work, it is desirable to also consider asymmetric influence between sponsors and cosponsors.
@@ 4.1 Accounting for the Accumulation and Decay of Influence with Time
We seek a cumulative-in-time notion of influence that incorporates decay with time. Intuitively, a legislator that has been in Congress for ten years has more ‘influence capital’ (i.e., social capital for the purpose of influence) than a legislator on their first day in Congress [47]. Additionally, a legislator that was very influential many years ago that has been inactive for a long time likely has experienced decay in their influence capital. To account for both of these ideas, we use an exponential decay function $e^{kt}$ so that accrued influence (from successful bills) decays as a function of the time $t$ since the introduction of a bill. The decay rate $k$ controls the half-life of a unit of influence between individuals that have successfully cosponsored a single passed bill. We examine influence half-lives of 6, 12, and 24 months by setting $k=\frac{\text{ln}(0.5)}{\text{half-life}}{\approx}\frac{-0.693}{\text{half-life}}$.
We consider an exponential influence decay for both $A[t,i,j]_{\text{pass}}$ (the number of bills that are introduced in the House at time $t$, include both representatives $i$ and $j$ as cosponsors, and eventually pass in the House) and $A[t,i,j]_{\text{tot}}$ (the number of bills, irrespective of their success, that are introduced in the House at time $t$ and include both representatives $i$ and $j$ as cosponsors). We then aggregate results month by month to yield tensors $C_{\text{pass}}$ and $C_{\text{tot}}$ with entries
@@
The entry $C[t,i,j]_{\text{pass}}$ is the number of bills that are cosponsored by both representatives $i$ and $j$ and pass in the House between months $1$ and $t$ (inclusive of months $1$ and $t$), with a separate decay applied each month. For example, we obtain the influence between representatives $i$ and $j$ in month 3 (while accounting for the influence that they accumulate in months $1$ and $2$) by calculating
@@
which incorporates decay each month. Analogously, the entry $C[t,i,j]_{\text{tot}}$ is the total number of bills (irrespective of their success) that are cosponsored by both representatives $i$ and $j$ between months $1$ and $t$, with decay each month.